# PIP-Net: Pedestrian Intention Prediction in the Wild

### v2

@@ III-C Cameras Features Integration
We define Sentinel Camera as a variable that indicates the index of the camera ($C_{ix}$) on which the target pedestrian has been observed. Using the Camera Index, we can adjust the pose and bounding box coordinates with respect to the sentinel camera. This task will be accomplished by the Shift unit, as shown in Figure 3 in cyan, which extends the global coordinates from the leftmost camera to the rightmost one, and applies these adjustments to the inputs. In this context, the Padding unit is responsible for generating a zero binary mask of size $c\times 512\times 512$, where the $c$-th dimension corresponds to the sentinel camera being set to one. Here, $c$ represents the number of cameras. As the aggregation module (A) shown in Figure 5, the binary masks are combined with the feature vectors generated by the VGG19 network for each $E_{sc}$ and $E_{cd}$. Subsequently, a pointwise convolution (Conv $1\times 1$) operator aggregates all the features across cameras. This process combines features from different channels (cameras) at each spatial location, allowing a weighted combination of input features.
@@ III-D Temporal and Attention Module
where $\sigma(.)$ denotes the logistic sigmoid function $x^{t}_{j}$ is the input feature at time step $t$. The reset and update gates at time step $t$ are denoted as $r^{t}_{j}$ and $z^{t}_{j}$, respectively, and the weights between the two units are represented by $W$. The hidden state at the previous time step and the current time step are represented by $h^{t-1}_{j}$ and $h^{t}_{j}$, respectively.
@@
In the tile of the network, the outputs of the attention modules are concatenated and then forwarded through the last attention module and an FC layer. The ultimate output, normalised to a range between zero and one using the Softmax function, represents the prediction for pedestrian crossing intention.
@@ IV Experiments
In this section, we conduct four distinct experiments to thoroughly assess the robustness of the proposed framework. Each experiment is designed to provide unique insights into various aspects of the model’s performance. We compare our model against the PIE dataset in four different intervals ranging from 1 to 4 seconds, allowing us to scrutinise the model’s predictive capabilities for an in-time response in different driving scenarios. The second experiment examines the impact of the introduced Categorical Depth and Local Motion features in enhancing the framework’s prediction accuracy. The third experiment evaluates the model’s generalisability and reliability on a diverse dataset from Waymo’s self-driving vehicles. This investigation ensures that the framework performs effectively across different real-world scenarios. Lastly, we investigate the scalability of the model by assessing the framework’s ability to handle one to three cameras simultaneously, expanding its view angles. This provides insights into the model’s efficiency when processing information from multiple cameras. This multi-faceted evaluation allows us to gain a detailed understanding of the model’s effectiveness, in comparison to a singular experiment.
@@ IV-A Datasets
Crossed Pedestrians 512 409
Not Crossed Pedestrians 1,328 1,072
IV-A1 PIE dataset
@@ IV-A2 Urban-PIP dataset
The dataset is recorded under various weather and daytime conditions in three geographical locations using a multi-sensor setup. This multi-modal dataset is collected via a combination of LiDAR, cameras, radar, and IMU sensors mounted on the ego-vehicle. LiDAR provides ${}^{\circ}$ field of view with approximately a 300-meter range by beaming out millions of laser pulses per second and measuring the time of laser beam flight from the sensor to the surface of an object, then reflecting from the object to the sensor on the ego-vehicle. The radar system has a continuous ${}^{\circ}$ view to track the presence and speed of road users in front, behind and sides of the vehicle. The front cameras (FL, F, and FR) simultaneously capture the traffic scene videos in HD format ($1920\times 1080$). The IMU module uses accelerometers and gyroscopes with input from GPS, maps, wheel speeds, as well as laser and radar measurements to provide position, velocity, and heading information to the vehicle.
@@
In this study, the experiments are conducted using camera sensors as they provide rich visual information, including detailed information about pedestrian behaviour, body language, and contextual information that can be crucial for predicting crossing intentions. Also, the affordability of camera sensors has made them a practical choice for current research on intention prediction. We annotated 1,481 pedestrian crossing intentions from the front cameras including 448 in the front-left camera, 541 from the front camera, and 492 from the front-right camera.
@@
To assess models limited to a single camera, we introduce a subset, named Frontal-Urban-PIP, focusing on pedestrians observed only by the front camera. This subset, featuring 55 pedestrians with crossing intentions and 129 without, ensures a fair comparison will be conducted with similar methods which are limited to a single camera only.
IV-B Implementation Settings
@@
The proposed model was executed on a CUDA parallel computing platform with an Nvidia Quadro RTX A6000 GPU and Intel Core i9 13900K 24-core processor and the Torch environment. PIP-Net was trained using the RMSProp optimiser. 256 hidden units were used for the GRUs, and the sigmoid ($\sigma$) activation function was applied to the GRUs for handling spatial kinematic data. To mitigate overfitting, a dropout rate of 0.5 was introduced after the attention block. Additionally, an L2 regularisation term of 0.0001 was incorporated into the last fully connected layer. A stride of 3 steps is used for each input sequence in the observation scene, resulting in a total of 10 frames for one second. This stride reduces frame redundancy and compensates for feature extraction time delay.
@@
PIP-Net-$\alpha$: This model is designed for a single camera setup. The model is trained on the PIE dataset and also evaluated against the PIE test set and Frontal-Urban-PIP datasets. The model doesn’t have the Camera Index pipeline within its architecture. Thereby, the outputs of the Max-pooling 3D (MP3D in Figure 3) are directly forwarded to the flattened block, and the Shift blocks are deactivated. A learning rate of $5\times 10^{-5}$ was used for 300 epochs with a batch size 10. The model was tested on the Frontal-Urban-PIP dataset to evaluate its generalisability. The input length stride was set to 1 because the dataset’s FPS was already equal to 10.
@@
TABLE II: SOTA Performance comparison on the PIE dataset.
@@
ATGC (2017) 0.59 0.55 0.39 0.33 0.47
Multi-RNN (2018) 0.83 0.80 0.71 0.69 0.73
SingleRNN (2020) 0.81 0.75 0.64 0.67 0.61
SFRNN (2020) 0.82 0.79 0.69 0.67 0.70
PCPA (2021) 0.87 0.86 0.77 0.75 0.79
CAPformer (2021) 0.88 0.80 0.71 0.69 0.74
PPCI (2022) 0.89 0.86 0.80 0.79 0.81
GraphPlus (2022) 0.89 0.90 0.81 0.83 0.79
MCIP (2022) 0.89 0.87 0.81 0.81 0.81
CIPF (2023) 0.91 0.89 0.84 0.85 0.83
PIP-Net-$\alpha$ (Ours) 0.91 0.90 0.84 0.85 0.84
IV-C Comparative Results
Table II highlights the performance of our method on the PIE dataset. The observation time ($m$) has been set to 16 frames, the same as the previous methods to ensure a fair comparison with other methods. PIP-Net-$\alpha$, achieves the highest values in all metrics, showcasing its prowess in capturing the crossing intention classification. These metrics include accuracy (Acc), precision, and recall rate, which quantify the model’s ability to accurately predict the binary classification task. Additionally, the area under the ROC curve (AUC), indicates the model’s proficiency in distinguishing between different classes, and F1 score, represents the harmonic mean of precision and recall rate.
In comparison to MultiRNN [24], SingleRNN [25], and SFRNN [23] models, which use a CNN encoder for visual features and RNN-based encoder-decoder structure, PIP-Net-$\alpha$ shows significant improvements by considering a new combination of input features. Transformer and graph-based architectures used in CAPformer [30] and GraphPlus [35] models have been less effective compared to PIP-Net with RNN-based architecture. The proposed feature fusion approach using seven input features, including three kinematic features and four contextual features, has also improved AUC by +1% compared to CIPF [11], which integrates eight distinct input features derived from pedestrians and vehicles through three fusion modules.
IV-C1 Crossing time prediction
@@
TABLE III: PIP-Net-$\alpha$ sub-variants using different input features on the PIE dataset. Abbreviations are: LM: Local Motion, GM: Global Motion, GD: Global Depth, and CD: Categorical Depth. The up arrow indicates an improvement and the down arrow indicates a decrement in comparison with the baseline variant.
Model GM LM GD CD Acc AUC F1
$\alpha^{0}$ 0.883 0.875 0.792
@@
$\alpha^{2}$ 0.889$\uparrow$ 0.887$\uparrow$ 0.801$\uparrow$
$\alpha^{3}$ 0.877$\downarrow$ 0.870$\downarrow$ 0.798$\uparrow$
$\alpha^{4}$ 0.904normal-$\uparrow$ 0.892normal-$\uparrow$ 0.829normal-$\uparrow$
$\alpha^{5}$ 0.875$\downarrow$ 0.871$\downarrow$ 0.789$\downarrow$
[…]
IV-C2 Features Importance
Recent studies [30, 10, 11, 12] have emphasised the importance and reliability of primary features including Bounding Box, Body Pose, Local Content, Vehicle Speed and Semantic Context, in their elaborate experiments. The baseline model, $\alpha^{0}$, comprises these primary features.
Initial experiments have demonstrated that excluding the input of the Bounding Box leads to 8.6% decrease in accuracy compared with the baseline. While omitting the Body Pose parameter only reduces accuracy by 3.5%. This represents less importance of body pose compared to the bounding box which may seem counter-intuitive. However, our further investigations confirm the bounding box data is notably more important and useful as it includes the pedestrian’s moving trajectory and tracking history over time, providing valuable spatio-temporal information. While body poses spatio-temporal information is not very important. The body pose only seems to matter when the pedestrian is about to cross the road in the last few frames before crossing. Otherwise, the body pose in previous moments, such as when the pedestrian is on the sidewalk, is redundant.
TABLE IV: The comparison table for the Frontal-Urban-PIP dataset.

### v3

@@ III-C Cameras Features Integration
We define Sentinel Camera as a variable that indicates the index of the camera ($C_{ix}$) on which the target pedestrian has been observed. Using the Camera Index, we can adjust the pose and bounding box coordinates with respect to the sentinel camera. This task will be accomplished by the Shift unit, as shown in Figure 3 in cyan, which extends the global coordinates from the leftmost camera to the rightmost one, and applies these adjustments to the inputs. In this context, the Padding unit is responsible for generating a zero binary mask of size $c\times 512\times 512$, where the $c$-th dimension corresponds to the sentinel camera being set to one. Here, $c$ represents the number of cameras. As the aggregation module (A) shown in Figure 5, the binary masks are combined with the input vectors i.e., $E_{sc}$ and $E_{cd}$. Subsequently, a pointwise convolution (Conv $1\times 1$) operator aggregates all the features across cameras. This process combines features from different channels (cameras) at each spatial location, allowing a weighted combination of input features.
@@ III-D Temporal and Attention Module
where $\sigma(\cdot)$ denotes the logistic sigmoid function $x^{t}_{j}$ is the input feature at time step $t$. The reset and update gates at time step $t$ are denoted as $r^{t}_{j}$ and $z^{t}_{j}$, respectively, and the weights between the two units are represented by $W$. The hidden state at the previous time step and the current time step are represented by $h^{t-1}_{j}$ and $h^{t}_{j}$, respectively.
@@
In the tile of the network, the outputs of the attention modules are concatenated and then forwarded through the last attention module and an FC layer. The ultimate output, normalised to a range between zero and one using the Sigmoid function, represents the predicted probability of a pedestrian crossing.
@@ IV Experiments
In this section, we conduct four distinct experiments to thoroughly assess the robustness of the proposed framework. Each experiment is designed to provide unique insights into various aspects of the model’s performance. We compare our model against the PIE dataset in four different intervals ranging from 1 to 4 seconds, allowing us to scrutinise the model’s predictive capabilities for an in-time response in different driving scenarios. Additionally, we evaluate the model’s performance under varying temporal resolution and observation time. The second experiment examines the impact of the introduced Categorical Depth and Local Motion features in enhancing the framework’s prediction accuracy. The third experiment evaluates the model generalisability and reliability on a diverse dataset from Waymo’s self-driving vehicles. This investigation ensures that the framework performs effectively across different real-world scenarios. Lastly, we investigate the scalability of the model by assessing the framework’s ability to handle one to three cameras simultaneously, expanding its view angles. This provides insights into the model’s efficiency when processing information from multiple cameras. This multi-faceted evaluation allows us to gain a detailed understanding of the model’s effectiveness in comparison to a singular experiment.
@@ IV-A Datasets
The evaluation metrics are based on [18], including accuracy (Acc), precision, and recall rate, which quantify the model’s ability to accurately predict the binary classification task. Additionally, the area under the ROC curve (AUC), indicates the model’s proficiency in distinguishing between different classes, and F1 score, represents the harmonic mean of precision and recall rate.
IV-A1 PIE dataset
@@
Lighting Conditions Daylight Daylight, Nighttime
Demographic Information Age, Gender None
@@
Crossed Pedestrians 519 409
Not Crossed Pedestrians 1,323 1,072
@@ IV-A2 Urban-PIP dataset
The dataset captures diverse traffic conditions, ranging from busy intersections to quieter residential streets, with dynamic environmental changes. Pedestrian behaviour may be influenced by the actions of nearby vehicles or other pedestrians, making intention prediction more challenging. It is recorded under various weather conditions, including sunny, cloudy, rainy, and foggy, in three geographical locations using a multi-sensor setup. This multi-modal dataset is collected via a combination of LiDAR, cameras, radar, and IMU sensors mounted on the ego-vehicle. LiDAR provides 360 field of view with approximately a 300-meter range by beaming out millions of laser pulses per second and measuring the time of laser beam flight from the sensor to the surface of an object, then reflecting from the object to the sensor on the ego vehicle. The radar system has a continuous 360 view to track the presence and speed of road users in front, behind and sides of the vehicle. The multi-camera perspective (left, front, and right) offers a richer contextual understanding of pedestrian interactions compared to single-camera datasets like PIE and JAAD. The cameras simultaneously capture the traffic scene videos in HD format ($1920\times 1080$), The IMU module uses accelerometers and gyroscopes with input from GPS, maps, wheel speeds, as well as laser and radar measurements to provide position, velocity, and heading information to the vehicle.
The Urban-PIP dataset captures both daylight and nighttime scenarios, making it more diverse and challenging than datasets such as PIE, JAAD, and STIP, which are limited to daylight conditions. The inclusion of low-light environments introduces additional challenges, where the visibility of pedestrians is reduced, and sensor data may become noisy, leading to challenges in detecting and tracking pedestrians’ behaviour accurately.
@@
In this study, the experiments are conducted using camera sensors as they provide rich visual information, including detailed information about pedestrian behaviour, their body language, and contextual information that can be crucial for predicting crossing intentions. Also, the affordability of camera sensors has made them a practical choice for current research on crossing prediction. We annotated 1,481 pedestrian crossing intentions from the front cameras including 448 in the front-left camera, 541 from the front camera, and 492 from the front-right camera.
IV-A3 Frontal Urban-PIP dataset
@@
To assess models limited to a single camera, we introduce a subset of the Urban-PIP dataset, named Frontal Urban-PIP (FU-PIP), focusing on pedestrians observed only by the front camera. This subset, featuring 55 pedestrians with crossing intentions and 129 without, ensures a fair comparison will be conducted with similar methods which are limited to a single camera only.
IV-B Hardware and Implementation Settings
All proposed models were executed on a CUDA parallel computing platform with an Nvidia Quadro RTX A5000 GPU, 64GB of RAM and an Intel Core i9 13900K 24-core processor and the Torch environment.
@@ IV-B Hardware and Implementation Settings
PIP-Net was trained using the RMSProp optimiser. 256 hidden units were used for the GRUs, and the sigmoid ($\sigma$) activation function was applied to the GRUs for handling spatial kinematic data. To mitigate overfitting, a dropout rate of 0.5 was introduced after the attention block. Additionally, an L2 regularisation term of 0.0001 was incorporated into the last fully connected layer.
@@
PIP-Net-$\alpha$: This model is designed for a single camera setup. The model is trained on the PIE dataset and evaluated against the PIE test set and the FU-PIP datasets. The model doesn’t have the Camera Index pipeline within its architecture. Thereby, the outputs of the Max-pooling 3D (MP3D in Figure 3) are directly forwarded to the flattened block, and the Shift blocks are deactivated. A learning rate of $5\times 10^{-5}$ was used for 300 epochs with a batch size 10. The model was tested on the FU-PIP dataset to evaluate its generalisability.
@@
IV-C Temporal Resolution and Observation Time
[…]
TABLE II: Performance results of PIP-Net-$\alpha$, memory usage, inference time, and observation time for different temporal strides and sequence lengths on PIE dataset.
s m OT Params M IT Acc AUC F1
1 10 0.3s 5.5M 1.4GB 55ms 0.913 0.907 0.835
2 0.5s 0.921 0.935 0.882
1 15 0.5s 5.9M 1.8GB 58ms 0.915 0.897 0.846
2 1.0s 0.907 0.891 0.836
1 20 0.6s 6.6M 2.5GB 62ms 0.908 0.911 0.853
2 1.3s 0.896 0.874 0.810
s: Temporal Stride; m: Sequence Length (frame); OT: Observation Time (second); Params: Number of model parameters (million); M: Memory usage (gigabyte); IT: Inference Time (millisecond).
To further investigate the impact of temporal dependencies, we conducted experiments by varying the observation sequence length $m$ while keeping the temporal stride fixed at 1 and 2. The objective was to determine whether increasing the number of frames to process improves prediction performance or introduces redundancy or irrelevant information. Table II presents the results, including the number of model parameters, memory usage, and inference time for different configurations. Our findings indicate that a shorter sequence length $(m=10)$ with a temporal stride of 2 $(s=2)$ achieves the highest prediction performance (AUC = 0.938, F1 = 0.882) while maintaining computational efficiency. This configuration $(s=2,m=10)$ keeps the observation time to 0.5 seconds like $(s=1,m=15)$ but with fewer frames processing. Furthermore, this setting requires fewer learnable parameters (5.5M) and lower memory usage (1.4GB), leading to a more efficient inference time of 55ms. These results suggest that increasing sequence length beyond a certain threshold may introduce redundant or irrelevant information about pedestrians’ crossing intentions.
TABLE III: Performance comparison on the PIE dataset.
@@ IV-C Temporal Resolution and Observation Time
ATGC [15] 0.59 0.55 0.39 0.33 0.47
Multi-RNN [30] 0.83 0.80 0.71 0.69 0.73
SingleRNN [31] 0.81 0.75 0.64 0.67 0.61
SFRNN [29] 0.84 0.82 0.72 0.75 0.80
PCPA [18] 0.87 0.86 0.77 0.75 0.79
CAPformer [23] 0.88 0.80 0.71 0.69 0.74
PPCI [43] 0.89 0.86 0.80 0.79 0.81
GraphPlus [49] 0.89 0.90 0.81 0.83 0.79
MCIP [11] 0.89 0.87 0.81 0.81 0.81
CIPF [12] 0.91 0.89 0.84 0.85 0.83
PIT [42] 0.91 0.92 0.82 0.84 0.81
VMIGI [44] 0.92 0.91 0.87 0.86 0.88
TrEP [39] 0.92 0.93 0.86 0.87 0.84
PIP-Net-$\alpha$ (Ours) 0.92 0.94 0.88 0.89 0.88
All models are evaluated at ETC = 0.5s. Results marked with are reported from the original article due to unavailable source code.
IV-D Comparative Results
Table III highlights the performance of our method on the PIE dataset. The observation time ($m$) has been set to 15 frames, the same as the previous methods, to ensure a fair comparison with other methods. PIP-Net-$\alpha$, achieves the highest values in all metrics, showcasing its prowess in capturing the crossing intention classification.
In comparison to MultiRNN, SingleRNN, and SFRNN, which rely on CNN encoders for visual features and RNN-based encoder-decoder structures, PIP-Net-$\alpha$ achieves significant improvements by integrating a more diverse set of input features, capturing both pedestrian kinematics and environmental context more effectively. Unlike transformer-based (CAPformer) and graph-based (GraphPlus) architectures, which struggle with motion generalisation and real-time inference, PIP-Net-$\alpha$ leverages an RNN-based architecture to maintain temporal continuity and robust sequential reasoning, resulting in higher predictive performance. Furthermore, the proposed feature fusion approach, combining three kinematic and four contextual features, enables a more comprehensive scene understanding, improving AUC by +4% compared to CIPF, which integrates eight pedestrian-vehicle interaction features.
While recent models such as PIT, VMIGI, and TrEP have introduced advanced techniques for long-range dependencies, relational learning, and uncertainty estimation, their effectiveness remains constrained by the inherent limitations of their respective architectures. PIT, incorporating self-attention mechanisms and Temporal Fusion Blocks, enhances sequence modelling but requires heavy computational resources, limiting real-time feasibility. VMIGI, which applies a multimodal fusion strategy with GCN-based relational learning, strengthens interaction reasoning but struggles with fast motion scenarios due to graph structure limitations. TrEP, adopting a transformer-based evidential prediction framework, explicitly models uncertainty by correlating it with human annotator disagreement, yet its reliance on pure transformer encoding leads to weaker sequential motion modelling. In contrast, PIP-Net-$\alpha$ effectively balances multimodal integration, temporal consistency, and computational efficiency, making it the most well-rounded solution for pedestrian crossing prediction in dynamic urban environments, modelling uncertainty and correlating it with human annotator disagreement.
IV-D1 Crossing Time Prediction
@@
IV-D2 Ablation Study
To investigate the inner workings of the proposed models, incorporating interpretability methods such as Grad-CAM [60] can enhance understanding by visualising which parts of the image influence predictions. However, in complex models that process heterogeneous input modalities and sequential data, these techniques may not always provide clear or actionable insights. Although VIMGI [44] successfully visualised the visual encoder, such methods often focus on individual components, missing feature interactions and kinematic data that affect predictions. Our approach, similar to studies such as CIPF [12], MCIP [11], TrEP [39] and CAPformer [23], emphasises performance-driven modelling, prioritising accuracy and efficiency. We have conducted comprehensive ablation studies to address interpretability concerns, providing valuable insights into our model’s decision-making processes.
The baseline model, $\alpha^{0}$, comprises the primary input features recognised in the task of pedestrian intention prediction, including Bounding Box, Body Pose, Local Content, Semantic Context, and Vehicle Speed [23, 11, 12, 13, 51]. Ablation experiments were conducted on the best-performing configuration, as presented in Table II, with s = 2 and m = 10. The results demonstrate that excluding the Bounding Box input leads to an 8.6% decrease in accuracy compared with the baseline, whereas omitting the Body Pose parameter results in a lower accuracy reduction of 3.5%. This represents less importance of body pose compared to the bounding box which may seem counter-intuitive. However, our further investigations confirm that the bounding box data is notably more important and useful as it includes the pedestrian’s moving trajectory and tracking history over time, providing valuable spatio-temporal information. While body poses spatio-temporal information is not very important. The body pose only seems to matter when the pedestrian is about to cross the road in the last few frames before crossing. Otherwise, the body pose in previous moments, such as when the pedestrian is on the sidewalk, is redundant.
The impact of removing Local Content causes a decrease in accuracy by 1.7%. It appears to lack comprehensive features on pedestrians’ intentions, given the wide variety in appearance and accessories they may have. Excluding Semantic Context leads to a 3.4% accuracy decrease as it includes details about road layout such as sidewalk positioning and drivable zones.
TABLE IV: PIP-Net-$\alpha$ variants using different input features.
Variant GM LM MD CD Acc AUC F1 IT
$\alpha^{0}$ 0.893 0.907 0.830 25
@@ IV-D2 Ablation Study
$\alpha^{2}$ 0.898$\uparrow$ 0.919$\uparrow$ 0.839$\uparrow$ +3
$\alpha^{3}$ 0.884$\downarrow$ 0.902$\downarrow$ 0.836$\uparrow$ +12
$\alpha^{4}$ 0.914$\uparrow$ 0.924$\uparrow$ 0.867$\uparrow$ +27
$\alpha^{5}$ 0.883$\downarrow$ 0.903$\downarrow$ 0.827$\downarrow$ +22
$\alpha$ 0.921$\uparrow$ 0.938$\uparrow$ 0.882$\uparrow$ +30
Features: GM is Global Motion, LM is Local Motion, MD is ManyDepth, and CD is Categorical Depth, IT: Inference Time (milliseconds). The up arrow indicates an improvement, and the down arrow indicates a decrement in comparison with the baseline variant.
Table IV shows a comparison of the two input features used in the proposed model, including scene motion and depth information, along with their impact on inference time. The $\alpha^{2}$ variant demonstrates that the local motion feature can exhibit superior performance when compared to the global motion feature (2d), which includes optical flow analysis of the entire scene. While optical flow is typically sensitive to any movement between consecutive frames, it appears that local motion can provide a coarse-grained feature and treat more concisely to account for pedestrians’ velocity and direction of movement, regardless of other irrelevant objects in the surroundings. Additionally, local motion introduces only a modest increase in inference time (+3ms), making it an efficient feature choice.
Regarding depth information features, the categorical depth feature proposed stands out as the most effective standalone feature, as evidenced by the results of variant $\alpha^{4}$ in Table IV, highlighting the significance of pedestrians’ distance to the ego-vehicle and their interactions with other road users in the traffic scene. However, this improvement comes at the cost of a notable increase in inference time (+27ms). Conversely, $\alpha^{3}$ performs the weakest among the variants when compared with the baseline, which utilises the global depth heatmap of the entire scene (2e). This underperformance may be attributed to the unstable estimation of depth for irrelevant surrounding objects. This is addressed in the categorical depth by focusing only on pedestrians and vehicles, and then applying per-instance normalisation (as shown in Figure 4).
Finally, the optimal outcome is attained by taking into account both the local motion and categorical depth map in $\alpha$, leading to enhancements in the baseline regarding Acc, AUC, and F1 score by 3.14%, 3.1%, and 6.2%, respectively. This improvement, however, comes with the highest inference time (+30ms), emphasising the trade-off between model complexity and real-time applicability. The full model ($\alpha$) achieves 55 milliseconds ($\approx$ 18 FPS) inference time while maintaining high prediction performance, ensuring its suitability for real-time autonomous systems.
IV-D3 Generalisation
The PIP-Net-$\alpha$ evaluation is presented against the SOTA methodologies on the FU-PIP dataset in Table V. Notably, the models have never seen the scenarios in their training phase. The outcomes of other methods were generated using the pre-trained weights they provided. Overall, models exhibit limited generalisation capability on Urban-PIP, suggesting that our dataset is a more challenging benchmark compared to PIE. Given its environmental variety and inclusion of nighttime scenes, we believe our findings are largely transferable to PIE and JAAD, which mainly comprise daylight conditions and may not fully capture real-world complexities.
Most models show improvements over ATGC across various metrics, with each demonstrating unique strengths. However, we witnessed a decrement in terms of performance for PCPA, PPCI, CAPformer, and GraphPlus. As far as the research curiosity demand, we investigated the architecture of these models, and it turned out they are suffering from low-quality global context (see Figure 8) and body pose features. This is caused by the feature extractor algorithms, i.e., the semantic segmentation and pose estimator they have used, which hinder the classifier from judging based on precise features.
In our model, we enhanced semantic context representation by employing panoptic segmentation instead of conventional semantic segmentation. Unlike semantic segmentation, which classifies each pixel into a category for a single image, panoptic segmentation processes video sequences to generate temporally consistent object segments, distinguishing between object instances and semantic categories simultaneously. However, failure cases arise in challenging scenarios, particularly in low-light conditions or occluded scenes. Figure 8 illustrates qualitative comparisons between Slot-VPS (used in our model) and DeepLabV3 [61] (used in previous works). During daytime, Slot-VPS [58] provides sharper instance boundaries and more stable object tracking. However, at night, we observe increased misclassification and temporal inconsistencies, where pedestrians may be incorrectly segmented as part of the background or falsely merged with nearby objects. These segmentation errors propagate to downstream predictions, potentially leading to incorrect pedestrian intention estimates, especially when multiple pedestrians are present.
TABLE V: Performance comparison on the Frontal Urban-PIP dataset.